# 2]The Blackett Laboratory, Prince Consort Road, Imperial College London, London SW7 2AZ, United Kingdom

### v1

@@ 2 Quantum Plasmonic Immunoassay Sensing
Figure 1: Quantum plasmonic immunoassay sensing. (a) Schematic diagram of the strong-coupling-enabled sensing device featuring a single analyte-emitter complex at the plasmonic hotspot. Key components are labelled. (b) Comparison of the sensing performance of the strong coupling (splitting-type) quantum sensor with classical (shifting-type) plasmonic sensors that feature dielectric and gold nanoparticle (AuNP) labels. (c) Photoluminescence (PL, blue solid line) and extinction (red dashed line) spectra for dimers with gap sizes of 2 – 6 nm and a fixed emitter transition dipole moment of 20 D. (d) Time-dependent polarization $P(t)$ of the single emitter label and corresponding envelopes (black solid lines), normalized in accordance with the incident field amplitude $E_{0}$ and plotted in a logarithmic representation. (e) Sensing figure of merit (FoM) in classical and quantum regimes as a function of the surface density of analyte-emitter complexes. The triangles indicate the FoM realized in each simulation with a random distribution of complexes, while the spheres indicate the average value for each surface density. Adapted with permission from Kongsuwan et al. [6]. Copyright 2019, American Chemical Society.
@@ 3 Near-field-driven Single-photon Emission and Multipartite Entanglement
Crucially, the transducer near-field can also be harnessed to generate dynamic bipartite entanglement, enabling one of the simplest two-qubit quantum logic gates, namely the controlled NOT (CNOT) gate. The degree of entanglement between the two qubits, whose energy levels are shown in Fig. 2a, can be systematically characterized via the concurrence $C$, with a maximum value of $C=1$ realized when, for instance, the system can be prepared in the Bell states $(|01\rangle\pm|10\rangle)/\sqrt{2}$ or $(|00\rangle\pm|11\rangle)/\sqrt{2}$ [17]. Fig. 2c shows the concurrence for two near-field-excited QDs in vacuum. Here, a single pulse from the NFT induces Rabi oscillations between the two-qubit ground state $|00\rangle$ and the state $|10\rangle$ featuring a single-qubit excitation. A second pulse from the NFT populates the state $|11\rangle$, after switching off the prepulse. Although $C$ remains much smaller than 1, it is similar to many proof-of-concept values reported using plasmonic nanocavities [13], and extends previous studies of plasmon-mediated entanglement into the strong coupling regime. Of note, the enhanced light-matter interaction facilitated by a near-zero-index host medium does not translate to improved entanglement between the qubits in this particular case, where the maximum achievable concurrence (about 0.07) is much smaller than for vacuum (about 0.2), in turn manifesting the need for an improved control over the coherences among the bipartite states.
This work has been further optimized and extended to the generation and dynamic modulation of three-qubit entanglement under ambient conditions [16]. As demonstrated for QDs embedded in silicon and multilayer graphene in Fig. 2d, repeated excitation of a tripartite Greenberger-Horne-Zeilinger state, $(|000\rangle+|111\rangle)/\sqrt{2}$, can be achieved with fidelities exceeding 0.5, and even approaching the maximum value of 1.0 at certain instants of time. Changes in the position of the QDs, or the choice of host medium and its thickness, may also alter the rate of oscillations and prove beneficial for optimizing a particular design. The ability to adeptly manipulate quantum superpositions and entanglement on nanometric spatial scales and ultrafast timescales will inevitably prove instrumental to the implementation of quantum networking concepts and error-correcting protocols for fault-tolerant QIP systems in chip-scale environments.
@@ 4 Outlook
The inherently small physical size of quantum plasmonic devices is itself an important quality. Their nanoscale compactness is naturally compatible with the drive for miniaturization in photonics and offers prospects for the development of chip-scale technologies that densely integrate many interconnected, quantum-optical components. This potential has been further enhanced by two complementary developments. Firstly, the emergence of plasmonic nanocircuitry [18, 19], which paves the way towards fully integrated, electrically-operable platforms for nanoplasmonic cavity and waveguide QED. Secondly, the recent proposal of dielectric engineering strategies for plasmonic hotspots and the control of plasmon-exciton strong coupling in low-dimensional semiconductors [14, 15], which is fully compatible with well-established, top-down lithographic fabrication approaches.

### v2

# Near-field Strong Coupling and Entanglement of Quantum Emitters for Room-temperature Quantum Technologies

@@ 2 Quantum Plasmonic Immunoassay Sensing
[…]
@@ 3 Near-field-driven Single-photon Emission and Multipartite Entanglement
Crucially, the transducer near-field can also be harnessed to generate dynamic bipartite entanglement, enabling one of the simplest two-qubit quantum logic gates, namely the controlled NOT (CNOT) gate. The degree of entanglement between the two qubits, whose energy levels are shown in Fig. 2a, can be systematically characterized via the concurrence $C$, with a maximum value of $C=1$ realized when, for instance, the system can be prepared in the Bell states $(|01\rangle\pm|10\rangle)/\sqrt{2}$ or $(|00\rangle\pm|11\rangle)/\sqrt{2}$ [17]. Fig. 2c shows the concurrence for two near-field-excited QDs in vacuum. Here, a single pulse from the NFT induces Rabi oscillations between the two-qubit ground state $|00\rangle$ and the state $|10\rangle$ featuring a single-qubit excitation. A second pulse from the NFT populates the state $|11\rangle$, after switching off the prepulse. Although $C$ remains much smaller than 1, it is similar to many proof-of-concept values reported using plasmonic nanocavities [13], and extends previous studies of plasmon-mediated entanglement into the strong coupling regime. Moreover, the achieved values of concurrence could potentially be improved via optimized NFT pulsing schemes and QD positioning in the near-field environment, or resonator designs based on alternative, low-loss plasmonic materials. Of note, the enhanced light-matter interaction facilitated by a near-zero-index host medium does not translate to improved entanglement between the qubits in this particular case, where the maximum achievable concurrence (about 0.07) is much smaller than for vacuum (about 0.2), in turn manifesting the need for an improved control over the coherences among the bipartite states.
This work has been further extended to the generation and dynamic modulation of three-qubit entanglement under ambient conditions [16]. As demonstrated for QDs embedded in silicon and multilayer graphene in Fig. 2d, repeated excitation of a tripartite Greenberger-Horne-Zeilinger state, $(|000\rangle+|111\rangle)/\sqrt{2}$, can be achieved with fidelities exceeding 0.5, and even approaching the maximum value of 1.0 at certain instants of time. Changes in the position of the QDs, or the choice of host medium and its thickness, may also alter the rate of oscillations and prove beneficial for optimizing a particular design. The ability to adeptly manipulate quantum superpositions and entanglement on nanometric spatial scales and ultrafast timescales will inevitably prove instrumental to the implementation of quantum networking concepts and error-correcting protocols for fault-tolerant QIP systems in chip-scale environments.
@@ 4 Outlook
The inherently small physical size of quantum plasmonic devices is itself an important quality. Their nanoscale compactness is naturally compatible with the drive for miniaturization in photonics and offers prospects for the development of chip-scale technologies that densely integrate many interconnected, quantum-optical components. This potential has been further enhanced by two complementary developments. Firstly, the emergence of plasmonic nanocircuitry [18, 19], which paves the way towards fully integrated, electrically-operable platforms for nanoplasmonic cavity and waveguide quantum electrodynamics. Secondly, the recent proposal of dielectric engineering strategies for plasmonic hotspots and the control of plasmon-exciton strong coupling in low-dimensional semiconductors [14, 15], which is fully compatible with well-established, top-down lithographic fabrication approaches.